# bHUN-REN Wigner Research Centre for Physics, Konkoly-Thege Miklos ut 29-33, 1121 Budapest, Hungary

### 09 Apr 2026 v2

@@ 2 Strong coupling expansion
$\mathcal{Z}^{(1,0)}(g)=e^{i\pi a}P^{(1,0)}\frac{(8\pi)^{2a}}{8\pi g}\left[1+((a-1)^{2}-\ell^{2})\sum_{k\geq 1}\frac{(-1)^{k}}{k!}\frac{f_{k}^{(1,0)}}{(8\pi g)^{k}}\right]$ (22)
@@
$\mathcal{Z}^{(0,1)}(g)=P^{(0,1)}\frac{(8\pi)^{-2a}}{8\pi g}\left[1+((a+1)^{2}-\ell^{2})\sum_{k\geq 1}\frac{(-1)^{k}}{k!}\frac{f_{k}^{(0,1)}}{(8\pi g)^{k}}\right]$ (27)
@@
$P^{(1,1)}=-\left(\frac{1}{2}-a\right)^{1+2a}\left(\frac{1}{2}+a\right)^{1-2a}$ (33)
@@ 3.1 $1/g$ expansion
$\mathcal{I}_{n}=\sum_{k=0}^{\infty}\left[(-1)^{n-1}\left(\frac{1}{(x_{k}^{+})^{n-1}}-\frac{1}{(y_{k})^{n-1}}\right)-\left(\frac{1}{(x_{k}^{-})^{n-1}}-\frac{1}{(y_{k})^{n-1}}\right)\right]$ (59)
@@
$\mathcal{I}_{n}\to\mathcal{I}^{(\{0\},\{\})}_{n}\equiv\mathcal{I}_{n}-(-1)^{n-1}\frac{2}{\left(x_{0}^{+}\right)^{n-1}}$ (60)
@@
$=\mathcal{I}_{n}-(-1)^{n-1}\sum_{l\in\delta^{+}}\frac{2}{\left(l+\frac{1}{2}-a\right)^{n-1}}+\sum_{j\in\delta^{-}}\frac{2}{(j+\frac{1}{2}+a)^{n-1}}$ (67)
@@ 3.2 Stokes constants
$S^{(\{0\},\{\})}$ $=\frac{(-1)^{\ell}}{2\pi i}\,\frac{\Gamma(1+\ell-a)}{\Gamma(\ell+a)}\frac{\Phi_{-}(-2i\pi x_{0}^{+})}{\partial_{x}\Phi_{+}(2i\pi x_{0}^{+})}(x_{0}^{+})^{2a-2}$
$S^{(\{\},\{0\})}$ $=\frac{(-1)^{\ell}}{2\pi i}\,\frac{\Gamma(1+\ell+a)}{\Gamma(\ell-a)}\frac{\Phi_{+}(-2i\pi x_{0}^{-})}{\partial_{x}\Phi_{-}(2i\pi x_{0}^{-})}(x_{0}^{-})^{-2a-2}$ (72)
@@
$S_{+}^{\{j\}}$ $=\frac{(-1)^{\ell}}{2\pi i}\,\frac{\Gamma(\ell-a)}{\Gamma(1+\ell+a)}\frac{\Phi_{-}(-2i\pi x_{j}^{+})}{\partial_{x}\Phi_{+}(2i\pi x_{j}^{+})}(x_{j}^{+})^{2a}$
$S_{-}^{\{j\}}$ $=\frac{(-1)^{\ell}}{2\pi i}\,\frac{\Gamma(\ell+a)}{\Gamma(1+\ell-a)}\frac{\Phi_{+}(-2i\pi x_{j}^{-})}{\partial_{x}\Phi_{-}(2i\pi x_{j}^{-})}(x_{j}^{-})^{-2a}$ (73)
@@
$S^{(\{0\},\{0\})}=-\left(\frac{1}{2}-a\right)^{1+2a}\left(\frac{1}{2}+a\right)^{1-2a}$
$S^{(\{0,1\},\{\})}=-\left(\frac{1}{2}-a\right)^{-1+2a}\left(\frac{3}{2}-a\right)^{-3+2a}\frac{\Gamma(\ell-a+1)\Gamma(\ell-a+2)}{\Gamma(\ell+a)\Gamma(\ell+a-1)}$
$S^{(\{0,1\},\{0\})}=\frac{(-1)^{\ell}}{4}\left(\frac{1}{2}+a\right)^{3-2a}\left(\frac{1}{2}-a\right)^{1+2a}\left(\frac{3}{2}-a\right)^{-1+2a}\frac{\Gamma(\ell-a+1)}{\Gamma(\ell+a)}$
$S^{(\{0,1,2\},\{\})}=\left(\frac{1}{2}-a\right)^{-1+2a}\left(\frac{3}{2}-a\right)^{-3+2a}\left(\frac{5}{2}-a\right)^{-5+2a}\frac{\Gamma(\ell-a+1)\Gamma(\ell-a+2)\Gamma(\ell-a+3)}{\Gamma(\ell+a)\Gamma(\ell+a-1)\Gamma(\ell+a-2)}$
$S^{(\{0,1\},\{0,1\})}=\frac{1}{144}\left(\frac{1}{2}-a\right)^{3+2a}\left(\frac{3}{2}-a\right)^{1+2a}\left(\frac{1}{2}+a\right)^{3-2a}\left(\frac{3}{2}+a\right)^{1-2a}$ (77)
@@
$\left.\frac{(-1)^{\ell}}{2\pi i}\,\frac{\Gamma(\ell-a)}{\Gamma(1+\ell+a)}\frac{\Phi_{-}(-2i\pi x_{k}^{+})}{\partial_{x}\Phi^{\{j\}}_{+}(2i\pi x_{k}^{+})}(x_{k}^{+})^{2a}\right|_{a\to a-2}=\left(1-\frac{x_{k}^{+}}{x_{j}^{-}}\right)^{2}\left(\left.S^{\{k\}}_{+}\right|_{a\to a-2}\right)\,.$ (80)
In case of $j=0$, $k=1$ and $j=1$, $k=0$, up to an overall minus sign, this reproduces the first lines of (3.2) and (3.2).
@@ 3.3 A practical example
$\mathcal{C}_{\ell}^{(4)}(g)=\left.S^{(\{0\},\{0\})}\mathcal{D}^{(\{0\},\{0\})}(g)\right|_{a=1/4}\,$, $\mathcal{C}_{\ell}^{(5)}(g)=e^{\frac{i\pi}{4}}(8\pi g)^{-\frac{1}{2}}\left.S^{(\{1\},\{\})}\mathcal{D}^{(\{1\},\{\})}(g)\right|_{a=1/4}$ (89)
@@
$\sum_{l\in\delta^{+}}\left(4l+1\right)+\sum_{j\in\delta^{-}}\left(4j+3\right)=n$ (90)
@@
$\mathcal{I}_{n}^{(\delta^{+},\delta^{-})}=\mathcal{I}_{n}-(-1)^{n-1}\sum_{l\in\delta^{+}}\frac{2}{\left(l+\frac{1}{4}\right)^{n-1}}+\sum_{j\in\delta^{-}}\frac{2}{(j+\frac{3}{4})^{n-1}}$ (94)
@@
$A_{\ell=1}(g)=\frac{e^{\frac{3\pi g}{4}}}{g^{\frac{21}{16}}}\Gamma\left(\frac{1}{4}\right)^{\frac{5}{4}}\Gamma\left(\frac{3}{4}\right)^{\frac{3}{4}}\frac{e^{\frac{1}{8}}}{2^{\frac{109}{48}}\pi^{\frac{21}{16}}A^{\frac{3}{2}}}$ (98)
@@
$\left.-\frac{1025(9\mathcal{I}_{2}+24)^{3}+44280(9\mathcal{I}_{2}+24)(9\mathcal{I}_{3}+32)+108000(27\mathcal{I}_{4}+128)}{1019215872\pi^{3}g^{3}}+\mathcal{O}\left(\frac{1}{g^{4}}\right)\right]$ (99)
@@
Finally, using (8), I pesent some analytic results for the strong coupling expansion of the cusp anomalous dimension.
@@
The perturbative part in (3.3) reproduces the expansion presented in Basso et al. (2008), while the non-perturbative corrections coincide with the results of Basso and Korchemsky (2009) and Dorigoni and Hatsuda (2015). With the method discussed above, any higher order non-perturbative contribution to the cusp anomalous dimension can be easily obtained in an extremely efficient way. The complete strong coupling structure of $\Gamma$ summarized in Bajnok et al. (2026a).
@@ 4.1 Stokes automorphism
$\ln\mathfrak{S}=\sum_{\omega}e^{-\tilde{g}\omega}\Delta_{\omega}$ (109)
@@ 4.3 Median resummation
$\mathcal{S}_{\mathrm{med}}=\mathcal{S}_{+}\circ\mathfrak{S}^{-1/2}=\mathcal{S}_{-}\circ\mathfrak{S}^{1/2}$ (125)
@@ Appendix A Bridge equations
$\mathcal{B}\left[\Phi^{(\delta)}\right](s)=\sum_{k\geq 0}\frac{\phi_{k}^{(\delta)}}{\Gamma(k+1)}s^{k}$ (131)
@@
$\mathfrak{S}=\exp\left(\sum_{\omega}e^{-\omega g}\Delta_{\omega}\right)\,$ (141)
@@
$\left(\Delta_{j}\right)^{2}=0,\qquad\left[\Delta_{j},\Delta_{l}\right]=0$ (147)

### 27 May 2026 v3

@@ 2 Strong coupling expansion
$\mathcal{Z}^{(1,0)}(g)=e^{i\pi a}P^{(1,0)}\frac{(8\pi)^{2a}}{8\pi g}\left[1+((a-1)^{2}-\ell^{2})\sum_{k\geq 1}\frac{(-1)^{k}}{k!}\frac{f_{k}^{(1,0)}}{(8\pi g)^{k}}\right]\,$, (22)
@@
$\mathcal{Z}^{(0,1)}(g)=P^{(0,1)}\frac{(8\pi)^{-2a}}{8\pi g}\left[1+((a+1)^{2}-\ell^{2})\sum_{k\geq 1}\frac{(-1)^{k}}{k!}\frac{f_{k}^{(0,1)}}{(8\pi g)^{k}}\right]\,.$ (27)
@@
$P^{(1,1)}=-\left(\frac{1}{2}-a\right)^{1+2a}\left(\frac{1}{2}+a\right)^{1-2a}\,.$ (33)
@@ 3.1 $1/g$ expansion
$\mathcal{I}_{n}=\sum_{k=0}^{\infty}\left[(-1)^{n-1}\left(\frac{1}{(x_{k}^{+})^{n-1}}-\frac{1}{(y_{k})^{n-1}}\right)-\left(\frac{1}{(x_{k}^{-})^{n-1}}-\frac{1}{(y_{k})^{n-1}}\right)\right]\,.$ (59)
@@
$\mathcal{I}_{n}\to\mathcal{I}^{(\{0\},\{\})}_{n}\equiv\mathcal{I}_{n}-(-1)^{n-1}\frac{2}{\left(x_{0}^{+}\right)^{n-1}}\,.$ (60)
@@
$=\mathcal{I}_{n}-(-1)^{n-1}\sum_{l\in\delta^{+}}\frac{2}{\left(l+\frac{1}{2}-a\right)^{n-1}}+\sum_{j\in\delta^{-}}\frac{2}{(j+\frac{1}{2}+a)^{n-1}}\,$, (67)
@@ 3.2 Stokes constants
$S^{(\{0\},\{\})}$ $=\frac{(-1)^{\ell}}{2\pi i}\,\frac{\Gamma(1+\ell-a)}{\Gamma(\ell+a)}\frac{\Phi_{-}(-2i\pi x_{0}^{+})}{\partial_{x}\Phi_{+}(2i\pi x_{0}^{+})}(x_{0}^{+})^{2a-2}\,$,
$S^{(\{\},\{0\})}$ $=\frac{(-1)^{\ell}}{2\pi i}\,\frac{\Gamma(1+\ell+a)}{\Gamma(\ell-a)}\frac{\Phi_{+}(-2i\pi x_{0}^{-})}{\partial_{x}\Phi_{-}(2i\pi x_{0}^{-})}(x_{0}^{-})^{-2a-2}\,.$ (72)
@@
$S_{+}^{\{j\}}$ $=\frac{(-1)^{\ell}}{2\pi i}\,\frac{\Gamma(\ell-a)}{\Gamma(1+\ell+a)}\frac{\Phi_{-}(-2i\pi x_{j}^{+})}{\partial_{x}\Phi_{+}(2i\pi x_{j}^{+})}(x_{j}^{+})^{2a}\,$,
$S_{-}^{\{j\}}$ $=\frac{(-1)^{\ell}}{2\pi i}\,\frac{\Gamma(\ell+a)}{\Gamma(1+\ell-a)}\frac{\Phi_{+}(-2i\pi x_{j}^{-})}{\partial_{x}\Phi_{-}(2i\pi x_{j}^{-})}(x_{j}^{-})^{-2a}\,.$ (73)
@@
$S^{(\{0\},\{0\})}=-\left(\frac{1}{2}-a\right)^{1+2a}\left(\frac{1}{2}+a\right)^{1-2a}\,$,
$S^{(\{0,1\},\{\})}=-\left(\frac{1}{2}-a\right)^{-1+2a}\left(\frac{3}{2}-a\right)^{-3+2a}\frac{\Gamma(\ell-a+1)\Gamma(\ell-a+2)}{\Gamma(\ell+a)\Gamma(\ell+a-1)}\,$,
$S^{(\{0,1\},\{0\})}=\frac{(-1)^{\ell}}{4}\left(\frac{1}{2}+a\right)^{3-2a}\left(\frac{1}{2}-a\right)^{1+2a}\left(\frac{3}{2}-a\right)^{-1+2a}\frac{\Gamma(\ell-a+1)}{\Gamma(\ell+a)}\,$,
$S^{(\{0,1,2\},\{\})}=\left(\frac{1}{2}-a\right)^{-1+2a}\left(\frac{3}{2}-a\right)^{-3+2a}\left(\frac{5}{2}-a\right)^{-5+2a}\frac{\Gamma(\ell-a+1)\Gamma(\ell-a+2)\Gamma(\ell-a+3)}{\Gamma(\ell+a)\Gamma(\ell+a-1)\Gamma(\ell+a-2)}\,$,
$S^{(\{0,1\},\{0,1\})}=\frac{1}{144}\left(\frac{1}{2}-a\right)^{3+2a}\left(\frac{3}{2}-a\right)^{1+2a}\left(\frac{1}{2}+a\right)^{3-2a}\left(\frac{3}{2}+a\right)^{1-2a}\,.$ (77)
@@
$\left.\frac{(-1)^{\ell}}{2\pi i}\,\frac{\Gamma(\ell-a)}{\Gamma(1+\ell+a)}\frac{\Phi_{-}(-2i\pi x_{k}^{+})}{\partial_{x}\Phi^{\{j\}}_{+}(2i\pi x_{k}^{+})}(x_{k}^{+})^{2a}\right|_{a\to a-2}=\left(1-\frac{x_{k}^{+}}{x_{j}^{+}}\right)^{2}\left(\left.S^{\{k\}}_{+}\right|_{a\to a-2}\right)\,.$ (80)
The differentiation in the denominator on the left-hand side acts on the new function $\Phi_{+}^{\{j\}}(x)$ defined in (66). Since $\Phi_{+}(x)$ in the numerator of (66) vanishes at $x=2i\pi x_{k}^{+}$ for every $k\geq 0$, therefore, the derivative in the denominator simplifies as:
$\partial_{x}\Phi^{\{j\}}_{+}(2i\pi x_{k}^{+})={\partial_{x}\Phi_{+}(2i\pi x_{k}^{+})\over\left(1-{x_{k}^{+}\over x_{j}^{+}}\right)^{2}}\,$, (81)
which results in the additional overall factor $\left(1-x_{k}^{+}/x_{j}^{+}\right)^{2}$ on the right-hand side of (80). In case of $j=0$, $k=1$ and $j=1$, $k=0$, up to an overall minus sign, this reproduces the first lines of (3.2) and (3.2).
@@ 3.3 A practical example
$\mathcal{C}_{\ell}^{(4)}(g)=\left.S^{(\{0\},\{0\})}\mathcal{D}^{(\{0\},\{0\})}(g)\right|_{a=1/4}\,$, $\mathcal{C}_{\ell}^{(5)}(g)=e^{\frac{i\pi}{4}}(8\pi g)^{-\frac{1}{2}}\left.S^{(\{1\},\{\})}\mathcal{D}^{(\{1\},\{\})}(g)\right|_{a=1/4}\,.$ (90)
@@
$\sum_{l\in\delta^{+}}\left(4l+1\right)+\sum_{j\in\delta^{-}}\left(4j+3\right)=n\,.$ (91)
@@
$\mathcal{I}_{n}^{(\delta^{+},\delta^{-})}=\mathcal{I}_{n}-(-1)^{n-1}\sum_{l\in\delta^{+}}\frac{2}{\left(l+\frac{1}{4}\right)^{n-1}}+\sum_{j\in\delta^{-}}\frac{2}{(j+\frac{3}{4})^{n-1}}\,.$ (95)
@@
$A_{\ell=1}(g)=\frac{e^{\frac{3\pi g}{4}}}{g^{\frac{21}{16}}}\Gamma\left(\frac{1}{4}\right)^{\frac{5}{4}}\Gamma\left(\frac{3}{4}\right)^{\frac{3}{4}}\frac{e^{\frac{1}{8}}}{2^{\frac{109}{48}}\pi^{\frac{21}{16}}A^{\frac{3}{2}}}\,.$ (99)
@@
$\left.-\frac{1025(9\mathcal{I}_{2}+24)^{3}+44280(9\mathcal{I}_{2}+24)(9\mathcal{I}_{3}+32)+108000(27\mathcal{I}_{4}+128)}{1019215872\pi^{3}g^{3}}+\mathcal{O}\left(\frac{1}{g^{4}}\right)\right]\,.$ (100)
@@
Finally, using (8), I present some analytic results for the strong coupling expansion of the cusp anomalous dimension.
@@
The perturbative part in (3.3) reproduces the expansion presented in Basso et al. (2008), while the non-perturbative corrections coincide with the results of Basso and Korchemsky (2009) and Dorigoni and Hatsuda (2015). With the method discussed above, any higher order non-perturbative contribution to the cusp anomalous dimension can be easily obtained in an extremely efficient way. The complete strong coupling structure of $\Gamma$ is summarized in Bajnok et al. (2026a).
@@ 4.1 Stokes automorphism
$\ln\mathfrak{S}=\sum_{\omega}e^{-\tilde{g}\omega}\Delta_{\omega}\,$, (110)
@@ 4.3 Median resummation
$\mathcal{S}_{\mathrm{med}}=\mathcal{S}_{+}\circ\mathfrak{S}^{-1/2}=\mathcal{S}_{-}\circ\mathfrak{S}^{1/2}\,.$ (126)
@@ Appendix A Bridge equations
$\mathcal{B}\left[\Phi^{(\delta)}\right](s)=\sum_{k\geq 0}\frac{\phi_{k}^{(\delta)}}{\Gamma(k+1)}s^{k}\,.$ (132)
@@
$\mathfrak{S}=\exp\left(\sum_{\omega}e^{-\omega g}\Delta_{\omega}\right)\,$, (142)
@@
$\left(\Delta_{j}\right)^{2}=0,\qquad\left[\Delta_{j},\Delta_{l}\right]=0\,$, (148)